OpenStudy (anonymous):

Consider the following expression: 8y^2+4y-2 ___________ 2y+1 Divide the polynomial in the numerator by the binomial in the denomiator

myininaya (myininaya):

4y ---------------- 2y+1| 8y^2+4y-2 -(8y^2+4y) --------------- -2

myininaya (myininaya):

\[\frac{8y^2+4y-2}{2y+1}=4y+\frac{-2}{2y+1}\]